OpenStudy (badsheep):

The average height of an apple tree is 4.00 meters. How long would it take an apple falling from that height to reach the ground? Given: g = -9.8 meters/second2

OpenStudy (hannahgrace4464):

.40 seconds i think

OpenStudy (badsheep):

0.90 1.0 1.53 2.3

OpenStudy (badsheep):

@hannahgrace4464

OpenStudy (hannahgrace4464):

.90 im pretty sure

OpenStudy (whpalmer4):

\[h = \frac{1}{2}gt^2\]you know \[h=4.00\]\[g=9.8\]solve for \(t\) and plug in the numbers.

OpenStudy (badsheep):

is t 4?

OpenStudy (whpalmer4):

\[h = \frac{1}{2}gt^2\]multiply both sides by 2\[2h = gt^2\]divide both sides by \(g\)\[\frac{2h}{g} = t^2\]take square root of both sides\[\sqrt{\frac{2h}{g}} = t = \sqrt{\frac{2*4}{9.8}} = \]

OpenStudy (badsheep):

.90?

OpenStudy (whpalmer4):

does 0.9 * 0.9 * (1/2)*9.8 approximately equal 4.0? If so, then yes.

OpenStudy (badsheep):

yeah thanks guys/girls

OpenStudy (whpalmer4):

by the way, you'll probably get faster help asking physics problems like this in the physics section rather than the chemistry section